1. If the velocity of a particle is given by v(t) = 4 + 4 and s(1) = 2, find the particular position function s(t).
Answer #1

Giren, v(t) = 46+4, SU)=2 Sit)= vet) de - Jệt+4). 462 1 / +46+C s(t) = 26²446+C SMN = 200 +461) + Co2 6+c=2 TC=-4 s(t)= 36?+4